OpenStudy (kramse18):

e^ydx-x(2xy+e^y)dy=0

OpenStudy (irishboy123):

\(e^ydx-x(2xy+e^y)dy=0\) exact or not?

OpenStudy (irishboy123):

nah

OpenStudy (kramse18):

its not

OpenStudy (irishboy123):

you should learn to typeset.

OpenStudy (pnathaniel56):

is anyone in here in a k12 school

OpenStudy (kramse18):

typeset?

OpenStudy (irishboy123):

\(t^{y^p} \dfrac{e}{s} e^t \)

OpenStudy (hughmungus14):

Hugh Mungus

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle e^ydx-(2x^2y+xe^y)dy=0}\) \(\color{black}{\displaystyle e^y\ne4xy+e^y\quad \Longrightarrow \quad\frac{dM}{dy}\ne\frac{dN}{dx}}\) try integrating factor.

OpenStudy (solomonzelman):

One of the two: \(\color{black}{\displaystyle g(x)=\frac{1}{N}[M_y-N_x]}\) \(\color{black}{\displaystyle h(y)=\frac{1}{M}[N_x-M_y]}\)

OpenStudy (sshayer):

\[e^ydx=(2x^2y+x e^y)dy\] \[e^y \frac{ dx }{ dy }=2x^2y+xe^y,\frac{ dx }{ dy }-x=2x^2y e ^{-y}\] divide by x^2 \[x^{-2}\frac{ dx }{ dy }-x^{-1}=2ye^{-y}\] Put~\[Put~x^{-1}=t\] \[-x^{-2}\frac{ dx }{ dy }=\frac{ dt }{ dy }\] \[-\frac{ dt }{ dy }-t=2ye^{-y},\frac{ dt }{ dy }+t=-2ye^{-y}\]

OpenStudy (sshayer):

\[I.F=e^{\int\limits 1dy}=e^y\] c.s. is \[t.e^y=-2 \int\limits ye^{-y}e^ydy=-2 \int\limits ydy+c=-2*\frac{ y^2 }{ 2 }+c=-y^2+c\] replace the value of t and get the result.